OpenStudy (anonymous):

Find the thirtieth term of the following sequence. -6, -4, -2, 0, ... 52 64 66

OpenStudy (tkhunny):

There is not enough information for a unique answer. You can make a broad, sweeping assumptions that the difference of +2 continues indefinitely, but the problem statement doesn't actually say that.

OpenStudy (anonymous):

I guess I either have to count it out or guess...? That's literally all it gave me.

OpenStudy (tkhunny):

It is common to encounter poorly written problems of this type. You do not QUITE have to "count it out". You can use your best algebra: Find the thirtieth term of the following sequence. -6, -4, -2, 0, ... Term #1: -6 Term #2: -4 = -6 + 2 Term #3: -2 = -6 + 2(2) Term #4: 0 = -6 + 3(2) Getting the idea? Term #15: -6 + 14(2) = ?? Term #23: -6 + 23(2) = ?? Term #30: Is what?